Foreign Exclusion and Tax Credit (LO 7.6)

Taxpayer L has income of $55,000 from Norway, which imposes a 40 percent income tax, and income of $45,000 from France, which imposes a 30 percent income tax. L has additional taxable income from U.S. sources of $200,000 and U.S. tax liability before credits of $105,000. What is the amount of the foreign tax credit? Do not round any division in your computations. If required, round your answer to the nearest dollar.

a. $35,500

b. $16,500

c. $45,000

d. $35,000

e. $100,000
Answer #1

Answer: (a) 35,500

Explanation

​​​​In the foreighn tax credit we include only those we earned in the foreighn country and subject to be related to foreighn income taxes in that the country. We can claimed and deduction or claim to credit on the foreighn country only.

calculations are as:

1. Norway tax liability =55000*40%=22000

2. France tax liability =45000*30%=13500

Total foreighn tax credit/paid =(22000+13500)=35,500

If we find the overall limitations value

So we apply this formula:

(net foreighn income/u.s taxable income)*u.s tax liability

​​​​​​Net foreighn income=55000+45000=100000

U.S taxable income=55000+45000+200000=300,000

U.S tax liability =105,000

=(100,000/300,000)*105000

35000